### v1

@@ Abstract
Generative models hold great potential, but only if one can trust the evaluation of the data they generate. We show that many commonly used quality scores for comparing two-dimensional distributions of synthetic vs. ground-truth data give better results than they should, a phenomenon we call the “grade inflation problem.” We show that the correlation score, Jaccard score, earth-mover’s score, and Kullback-Leibler (relative-entropy) score all suffer grade inflation. We propose that any score that values all datapoints equally, as these do, will also exhibit grade inflation; we refer to such scores as “equipoint” scores. We introduce the concept of “equidensity” scores, and present the Eden score, to our knowledge the first example of such a score. We found that Eden avoids grade inflation and agrees better with human perception of goodness-of-fit than the equipoint scores above. We propose that any reasonable equidensity score will avoid grade inflation. We identify a connection between equidensity scores and Rényi entropy of negative order. We conclude that equidensity scores are likely to outperform equipoint scores for generative models, and for comparing low-dimensional distributions more generally.
@@ I Introduction
The ability to compare pairs of two-dimensional distributions robustly and accurately is critical in machine learning, and in data analysis more generally. This ability is especially valuable in the context of generative models [1] to measure model quality by assessing how well synthetic data fits the training data. For example, generative modeling of tabular data (where each row is a datapoint and each column is a feature) has important uses across many fields [2, 3, 4, 5]. Generative modeling is of great interest for imaging and multi-modal data as well [1, 6, 7]. Although datasets for the latter applications are generally high-dimensional, two-dimensional analysis plays a critical role in quality assessment, either for testing the fidelity of pairwise relationships or via dimensionality reduction techniques such as PCA, tSNE, or UMAP [8, 9, 10].
@@
Despite its ubiquity, the task of comparing two-dimensional distributions is non-trivial and has led to much work in developing quality scores. At a high level, quality scores can be classified as statistical vs. functional. Statistical scores are designed to demonstrate that some statistic, for example the mean of one of the features, has the same value in the synthetic dataset as in the real dataset (up to sampling error). In contrast, functional scores are meant to show that the outcome of some procedure, for example a classifier’s inference or human inspection, is indistinguishable regardless of whether the input data is real or synthetic (again, up to sampling error) [11]. Statistical scores have the advantage of being easier to calculate and of being generalizable from dataset to dataset; functional scores often take more human and/or compute time and are more likely to be idiosyncratic to a particular dataset. For these reasons, statistical scores are of general interest.
@@
Statistical scores appear often in the literature on generative modeling, especially of synthetic tabular data. For example, two widely-used and easy-to-compute statistics are the correlation coefficient $R$ and its square $R^{2}$ (the coefficient of determination), which measure the joint distribution between pairs of features (columns). In turn, two common types of correlation coefficient are Pearson’s, which measures linear relationships, and Spearman’s, which is a generalization for any monotonic relationship. For a given pair of columns, one can calculate Pearson’s or Spearman’s $R$ for the real data ($R_{p}$), do the same for the synthetic data ($R_{q}$), and use them to calculate a correlation score, for example as $21-|R_{p}-R_{q}|/$ 2 [12] or $1-|R^{2}_{p}-R^{2}_{q}|$. (Division normalizes the range to 0-to-1; subtraction converts the difference into a similarity score.) These two versions correlate closely with each other; the first is used more often in the literature, where it is also known as the correlation similarity.
@@
There are two important challenges to using the correlation score as a measure of fit between a pair of two-dimensional distributions. The first is the well-known problem that many different distributions can have the same correlation coefficient, as illustrated by examples such as Anscombe’s quartet and the Datasaurus Dozen [13, 14] (Fig. 1). Second, there are many more ways to get a low correlation than a high correlation. These two challenges can have the following effect. Suppose two columns in the real/training data have some non-random relationship that happens to have a near-zero correlation, for example like columns in the Datasaurus Dozen, and that the synthetic/generated data fails to learn this relationship, resulting in a random distribution. Because random data has zero correlation (up to sampling variance), the correlation score between the real and synthetic data will be high, despite the low quality of the synthetic data. In principle such issues can to lead to inappropriately high quality scores, a “grade inflation” problem that can make a generative model look better than it is.
@@
To further characterize the grade inflation problem, here we evaluate several additional quality scores: the earth-mover’s score, Jaccard score, and Kullback-Leibler or relative-entropy score. We find that all can fall prey to grade inflation. To address this situation, we introduce and describe a new score, the Eden score, that appears to avoid grade inflation. Our investigation focuses on two-dimensional distributions and pairwise relationships, which are often effective—sometimes “unreasonably” so [15, 16]—characterizing high-dimensional data.
@@ II-A Datasets
We generated synthetic data from both in-house two-dimensional toy datasets (“Dart’,’ “Trimodal," and “Stripes”) and the following high-dimensional machine-learning datasets obtained from the University of California Irvine machine-learning repository (UCIMLR): “Covertype,” “Communities and Crime,” “Energy Efficiency,” and “Rice (Cammeo and Osmancik)” (UCIML ID nos. 31, 183, 242, and 545, respectively). The latter were retrieved using the lucie Python package [2]. Anscombe’s Quartet was obtained via the seaborn Python package. The Datasaurus Dozen were obtained via https://www.autodesk.com/content/dam/autodesk/www/autodesk-reasearch/Publications/pdf/SameStatsCode.zip from the file named DatasaurusDozen.tsv.
@@ II-B Generative models and KDEs
II-C Quality scores
We consider five different quality scores. The correlation score, the earth mover’s score, the Jaccard score, and the Kullback-Leibler (KL) divergence score are based on previous work [24]; the Eden score is newly described below. Correlation score was defined using Pearson’s $R$ as $21-|R_{p}-R_{q}|/$ 2, with $p$ and $q$ the two distributions (e.g. real vs. synthetic data), reflecting the most common definition we found the literature. The earth-mover’s score was defined as $e^{-k\cdot\mathrm{EMD}}$, where $k$ is a scaling factor that can be chosen to scale the score as desired (here, $k=1$), and $\mathrm{EMD}$ is the normalized earth-mover’s distance from the pyemd python package [25]. Details for the remaining scores follow.
II-D Jaccard score implementation
@@
The size of the union is straightforward to define: it is the total number of points. In contrast, the intersection is a fuzzy notion due to points in the real and synthetic data generally taking continuous values and therefore generally being disjoint sets. Therefore, to define the intersection, we fit one Gaussian KDE model to the real data and a second Gaussian KDE to the synthetic data. We denote these real and synthetic KDEs $\hat{f}_{p}$ and $\hat{f}_{q}$, respectively. We take the intersection to be:
@@ II-F Eden score
In the Eden score, the difference between two KDEs is calculated by calculating the difference for each successive ring (or “annulus”) defined by the contours of the KDEs, and averaging these differences. For each annulus in the real/synthetic data, we define a per-annulus similarity score. Without loss of generality, label the outermost annulus 0, the second-outermost annulus 1, and so on. (The unbounded region outside of all contours is excluded.) Define the $i^{th}$ per-annulus score by a variation of the Jaccard formula (Eq. 1), with the area playing the role of size:
@@
Here $n_{\text{annuli}}$ is the number of annuli (note: for simplicity we refer to the innermost contour an annulus, even though technically it encloses a disk [possibly more than one]). $n_{\text{annuli}}=5$ was used.
To estimate the areas in the numerator and denominator of Eq. 9, a Monte-Carlo method is used: a bounding rectangle centered at the data is sprinkled with a large number of uniformly distributed points; the number of points that lie inside the union/intersection of a real and a synthetic annulus $i$ are counted. To determine whether a point lies inside a union/intersection of two annuli, we compute the likelihood of the point under $\hat{f}_{r}$ and $\hat{f}_{s}$ and check whether the likelihood falls within the likelihood range that defines two adjacent contours. Note the (minor) stochasticity due to KDE sampling and the Monte-Carlo procedure.
@@ II-G Validation
We used human visual inspection as a gold standard for comparison of two-dimensional distributions. 39 pairs of plots were shown to 20 human raters. Each plot consisted of two superimposed KDEs in different colors, corresponding to training data and synthetic data output by a generative model. Plots were chosen so that the plot with the higher Eden score received a lower score according to at least one of the other score. These results were not shared with the human raters, making this a blind test. Each human rater was asked to choose the plot in which the contours matched better, considering all contours; the interpretation of “better” was otherwise left up to the rater. The 39 pairs corresponded to 3 repeats of each of 13 unique pairs of columns. For each repeat, the pair was subjected to rotations or color swaps, and the left-right order of the plots was randomized; this allowed for a per-person test of consistency (to be able to assess whether the same plot picked all three times regardless of position, orientation, and color, or not). Raters all had a background in data science, science, and/or engineering, to increase the likelihood of exposure to/familiarity with the general practice of data presented as KDEs. Separately (i.e. without the rater), the fit in each plot was scored according to each scoring method (correlation, earth-mover’s, Jaccard, KL, and Eden) to determine the higher-scoring plot in each pair. Cohen’s kappa was used to calculate the agreement between each rater and each scoring method. Mann-Whitney U test was used to test the null hypotheses that distributions are the same location (Python, scipy.stats.mannwhitneyu).
@@ III-A Correlation score
To illustrate how the correlation score can lead to grade inflation, we first measured this score between pairs of distributions from Anscombe’s quartet. The four distributions in this set were devised to show how different distributions can have identical Pearson’s $R$ (to several digits). Fig. 1a illustrates the outcome: a perfect correlation score, despite the two distributions differing materially from each other. This is grade inflation. The “dino” dataset from the Datasaurus Dozen illustrates the compounding problem that arises as a result of low-correlation distributions being more common than high-correlation ones (Fig. 1b). Specifically, treating dino as the training set, we randomly intialized a generative model with the same x and y means and standard deviations, and sampled from that model—without any training. The dino dataset has a Pearson’s $R$ of close to zero. Random data also has a Pearson’s $R$ of close to zero. Because these values are similar, the correlation score is nearly perfect—0.97 in the sample in Fig. 1b—despite the model having learned literally nothing beyond the location and scale of the data. Again, this is grade inflation.
To illustrate the phenomenon on datasets and generative models used in real-world data science, we applied a variety of models to a selection of datasets from the UCIMLR and measured the correlation score between the synthetic/generated data and the real/training data, for representative pairs of columns (Fig. 2a-d). Even when fits were low quality by eye, the correlation score was universally high, ranging from 0.903 to 0.994 (Table I), demonstrating grade inflation. As a positive control, we also scored a high-quality model fit of an in-house dataset called Dart (Fig. 2e). Not surprisingly, the correlation score was also excellent, at 0.981, but notably this score was actually lower than the 0.994 achieved by the low-quality fit in Fig. 2a (Table I). Thus, correlation score has difficulty differentiating between high- and low-quality fits, leading to grade inflation for some poor fits.
Figure 2: Fits scored in Table I. Left to right: real data (blue), generated synthetic data (yellow), and KDEs. a-d are considered low-quality fits; e is considered a high-quality fit. Datasets, sizes, features, and models are as labeled. UCIMLR=UC Irvine Machine-Learning Repository. GC=Gaussian copula. nFlow=normalizing flow. ICA=independent component analysis.
@@ III-B Earth-mover’s score
A more sophisticated score derives from the earth-mover’s distance (EMD; a.k.a. the Mallow, Wasserstein, or Kantorovich-Rubenstein distance) [27, 28, 29, 30], which is the basis of the Frechet inception distance that is commonly used in machine learning [31]. If each two-dimensional distribution is a pile of sand, the EMD is the minimum amount of work required to transform one distribution into the other by moving sand around. EMD is easily converted to a score with range $(0,1)$ via exponentiation (Methods). To apply this score, the finite collections of datapoints in each distribution binned into histograms (the algorithm in [25] does not require smoothing the data by KDE, though that could have been done as well). In its native form, the EMD has the undesirable feature of being sensitive to the overall number of data points; to avoid this scale-dependence, the real and synthetic data are generally normalized, as we do here, before computing the EMS.
We calculated the earth-mover’s score for each fit in Fig. 2 and found this score also exhibits grade inflation (Table I). Specifically, it resulted in high scores in all cases where the training data exhibited sharp boundaries, regardless of whether the fit was low-quality, as in Figs. 2b and d (scores of 0.982 and 0.959, respectively), or high-quality, as in Fig. 2e (a score of 0.949). The earth-mover’s score did, however, result in (appropriately) low scores in the low-quality fits of the two multimodal distributions, Figs. 2a and c (scores of 0.124 and 0.379, respectively). Of note, scaling the scores by choosing a different value of $k$ does not resolve the grade inflation in Figs. 2b and d, because scaling so that low-quality fits have lower scores has the adverse effect of also reducing the score of high-quality fits; because changing $k$ does not change the ordering of scores, $k$ cannot thread this needle. Also of note, the earth-mover’s score can be highly sensitive to outliers, because the optimal transport might be such that the outliers have to be moved over further distances compared to more central points (see Discussion).
TABLE I: Scores for the fits in Fig. 2. Fig. 2a-d are considered low-quality fits deserving low scores; Fig. 2e is considered a high-quality fit deserving high scores. See Results for details. Standard deviations are reported for scores whose calculation involves a stochastic element (Methods).

### v2

@@ Abstract
Generative models hold great potential, but only if one can trust the evaluation of the data they generate. We show that many commonly used quality scores for comparing two-dimensional distributions of synthetic vs. ground-truth data give better results than they should, a phenomenon we call the “grade inflation problem.” We show that the correlation score, Jaccard score, earth-mover’s score, and Kullback-Leibler (relative-entropy) score all suffer grade inflation. We propose that any score that values all datapoints equally, as these do, will also exhibit grade inflation; we refer to such scores as “equipoint” scores. We introduce the concept of “equidensity” scores, and present the Eden score, to our knowledge the first example of such a score. We find that Eden avoids grade inflation and agrees better with human perception of goodness-of-fit than the equipoint scores above. We propose that any reasonable equidensity score will avoid grade inflation. We identify a connection between equidensity scores and Rényi entropy of negative order. We conclude that equidensity scores are likely to outperform equipoint scores for generative models, and for comparing low-dimensional distributions more generally.
@@ I Introduction
The ability to compare pairs of two-dimensional distributions robustly and accurately is critical in machine learning, and in data analysis more generally. This ability is especially valuable in the context of generative models [1] to measure model quality by assessing how well synthetic data fits training data. For example, generative modeling of tabular data, where each row is a datapoint and each column is a feature, has important uses across many fields [2, 3, 4, 5]. Generative modeling is of great interest for imaging and multi-modal data as well [1, 6, 7]. Although datasets for the latter applications are generally high-dimensional, two-dimensional analysis plays a critical role in quality assessment, either for testing the fidelity of pairwise relationships or via dimensionality reduction techniques such as PCA, tSNE, or UMAP [8, 9, 10].
@@
Despite its ubiquity, the task of comparing two-dimensional distributions is non-trivial and has led to much work in developing quality scores. At a high level, quality scores can be classified as statistical vs. functional. Statistical scores are designed to demonstrate that some statistic, for example the mean of one of the features, has the same value in the synthetic dataset as in the real dataset (up to sampling error). In contrast, functional scores are meant to show that the outcome of some procedure, for example inference, is indistinguishable regardless of whether the input data is real or synthetic (again, up to sampling error) [11]. Statistical scores have the advantage of being easier to calculate and of being generalizable from dataset to dataset; functional scores often take more human and/or compute time and are more likely to be idiosyncratic to a particular dataset. For these reasons, statistical scores are of general interest.
@@
Statistical scores appear often in the literature on generative modeling, especially of synthetic tabular data. For example, two widely-used and easy-to-compute statistics are the correlation coefficient $R$ and its square $R^{2}$ (the coefficient of determination), which measure the joint distribution between pairs of features (e.g. columns in tabular data). In turn, two common types of correlation coefficient are Pearson’s, which measures linear relationships, and Spearman’s, which is a generalization for any monotonic relationship. For a given pair of columns, one can calculate Pearson’s or Spearman’s $R$ for the real data ($R_{p}$), do the same for the synthetic data ($R_{q}$), and use them to calculate a correlation score, for example as $21-|R_{p}-R_{q}|/$ 2 [12] or $1-|R^{2}_{p}-R^{2}_{q}|$. (Division normalizes the range to 0-to-1; subtraction converts the difference into a similarity score.) These two versions correlate closely with each other; the first is used more often in the literature, where it is also known as the correlation similarity.
@@
There are two important challenges to using the correlation score as a measure of fit between a pair of two-dimensional distributions. The first is the well-known problem that many different distributions can have the same correlation coefficient, as illustrated by examples such as Anscombe’s quartet [13, 14] (Fig. 1a). Second, there are many more ways to get a low correlation than a high correlation. These two challenges can have the following effect. Suppose two columns in the real/training data have some non-random relationship that happens to have a near-zero correlation, for example like columns in the Datasaurus Dozen datasets, and that the synthetic/generated data fails to learn this relationship, resulting in a random distribution (Fig. 1b). Because random data has zero correlation (up to sampling variance), the correlation score between the real and synthetic data will be high, despite the low quality of the synthetic data. In principle such issues can to lead to inappropriately high quality scores, a “grade inflation” problem that can make a generative model look better than it is.
To further characterize the grade inflation problem, here we evaluate several additional quality scores: the earth-mover’s score, Jaccard score, and Kullback-Leibler or relative-entropy score. We find that all can fall prey to grade inflation. To address this problem, we introduce and describe a new score, the Eden score, that appears to avoid grade inflation. Our investigation focuses on two-dimensional distributions and pairwise relationships, which are often effective—sometimes “unreasonably” so [15, 16]—at capturing key relationships in high-dimensional data.
@@ II-A Datasets
We generated synthetic data from both in-house two-dimensional toy datasets (called Dart, Trimodal, and Stripes) and the following high-dimensional machine-learning datasets obtained from the University of California Irvine machine-learning repository (UCIMLR): Covertype, Communities and Crime, Energy Efficiency, and Rice (Cammeo and Osmancik) (UCIML ID nos. 31, 183, 242, and 545, respectively). The latter were retrieved using the lucie Python package [2]. Anscombe’s Quartet was obtained via the seaborn Python package. The Datasaurus Dozen were obtained via https://www.autodesk.com/content/dam/autodesk/www/autodesk-reasearch/Publications/pdf/SameStatsCode.zip from the file DatasaurusDozen.tsv.
@@ II-B Generative models and KDEs
II-C Correlation and earth-mover’s scores
We consider five different quality scores. The correlation score, the earth mover’s score, the Jaccard score, and the Kullback-Leibler (KL) divergence score are based on previous work [24]; the Eden score is newly described in this work (see below). Correlation score is defined using Pearson’s $R$ as $21-|R_{p}-R_{q}|/$ 2, with $p$ and $q$ the two distributions (e.g. real vs. synthetic data), reflecting the most common definition we found the literature. The earth-mover’s score is defined as $e^{-k\cdot\mathrm{EMD}}$, where $k$ is a scaling factor that can be chosen to scale the score as desired (here, $k=1$), and EMD is the normalized earth-mover’s distance from the pyemd python package [25]. Details for the remaining scores follow.
II-D Jaccard score
@@
The size of the union is straightforward to define: it is the total number of points. In contrast, the intersection is a fuzzy notion due to points in the real and synthetic data generally being disjoint sets. Therefore, to define the intersection, we fit one Gaussian KDE model to the real data and a second Gaussian KDE to the synthetic data. We denote these real and synthetic KDEs $\hat{f}_{p}$ and $\hat{f}_{q}$, respectively. We take the intersection and union to be:
@@ II-F Eden score
In the Eden score, the difference between two KDEs is calculated by calculating the difference for each successive ring or “annulus” defined by the contours of the KDEs, and averaging these differences. For each annulus in the real/synthetic data, we define a per-annulus similarity score. Without loss of generality, label the outermost annulus 0, the second-outermost annulus 1, and so on. (The unbounded region outside of all contours is excluded.) Define the $i^{th}$ per-annulus score by a variation of the Jaccard formula (Eq. 1), with the area playing the role of size:
@@
Here $n_{\text{annuli}}$ is the number of annuli. Note, technically the innermost annulus is usually a disk, and any annulus can consist of non-contiguous densities (e.g. multiple peaks). $n_{\text{annuli}}=5$ was used.
To estimate the areas in the numerator and denominator of Eq. 9, a Monte-Carlo method is used: a bounding rectangle centered at the data is sprinkled with a large number of uniformly distributed points; the number of points that lie inside the union/intersection of a real and a synthetic annulus $i$ is counted. To determine whether a point lies inside a union/intersection of two annuli, we compute the likelihood of the point under $\hat{f}_{r}$ and $\hat{f}_{s}$ and check whether the likelihood falls within the likelihood range that defines two adjacent contours. As with the KL score, this process introduces minor stochasticity.
@@ II-G Validation
We used human visual inspection as a gold standard for comparison of two-dimensional distributions. 39 pairs of plots were shown to 20 human raters. Raters all had a background in data science, science, and/or engineering, to increase the likelihood of exposure to/familiarity with the general practice of data presented as KDEs. Each plot consisted of two superimposed KDEs in different colors, corresponding to training data and synthetic data output by a generative model. Plots were chosen so that the plot with the higher Eden score received a lower score according to at least one of the other scores. Scores were not shared with the human raters, making this a blind test.
Each human rater was asked to choose the plot in which the contours matched better, considering all contours; the interpretation of “better” was otherwise left up to the rater. The 39 pairs corresponded to 3 repeats of each of 13 unique pairs of columns. For each repeat, the pair was subjected to rotations or color swaps, and the left-right order of the plots was randomized; this allowed for a per-person test of consistency, to assess whether the rater picked the same plot all three times regardless of position, orientation, and color. Separately (i.e. without the rater), the fit in each plot was scored according to each scoring method—correlation, earth-mover’s, Jaccard, KL, and Eden—to determine the higher-scoring plot in each pair. Cohen’s kappa was used to calculate the agreement between each rater and each scoring method. The Mann-Whitney U test was used to test the null hypotheses that each two scoring methods were equally good (Python, scipy.stats.mannwhitneyu).
@@ III-A Correlation score
To illustrate how the correlation score can lead to grade inflation, we first measured this score between pairs of distributions from Anscombe’s quartet. The four distributions in this set were devised in the 1973 to show how different distributions can have identical Pearson’s $R$ (to several digits). Fig. 1a illustrates the outcome: a perfect correlation score, despite the two distributions differing materially from each other. This is grade inflation. The “dino” dataset from the Datasaurus Dozen illustrates the compounding problem that arises as a result of low-correlation distributions being more common than high-correlation ones (Fig. 1b). Specifically, treating dino as the training set, we randomly intialized a generative model with the same x and y means and standard deviations, and sampled from that model without any training. The dino dataset has a Pearson’s $R$ of close to zero. Random data also has a Pearson’s $R$ of close to zero. Because these values are similar, the correlation score is nearly perfect—0.97 in the sample in Fig. 1b—despite the model having learned literally nothing beyond the location and scale of the data. Again, this is grade inflation.
To illustrate the phenomenon on datasets and generative models used in real-world data science, we applied a variety of models to a selection of datasets from the UCIMLR and measured the correlation score between the synthetic/generated data and the real/training data, for representative pairs of columns (Fig. 2a-d). Even when fits were low quality by eye, the correlation score was universally high, ranging from 0.903 to 0.994 (Table I), demonstrating grade inflation. As a positive control, we also scored a high-quality model fit of an in-house dataset called Dart (Fig. 2e). Not surprisingly, the correlation score was also excellent, at 0.981, but notably this score was actually lower than the 0.994 achieved by the low-quality fit in Fig. 2a (Table I). Thus, correlation score has difficulty differentiating between high- and low-quality fits, leading to grade inflation for some of the latter.
Figure 2: Fits scored in Table I. Left to right: real data (blue), generated synthetic data (yellow), and KDEs. a is considered a very low-quality fit; b-d are considered low-quality fits; e is considered a high-quality fit. Datasets, sizes, features, and models are as labeled. UCIMLR=UC Irvine Machine-Learning Repository. GC=Gaussian copula. nFlow=normalizing flow. ICA=independent component analysis.
@@ III-B Earth-mover’s score
A more sophisticated score derives from the earth-mover’s distance (EMD; a.k.a. the Mallow, Wasserstein, or Kantorovich-Rubenstein distance) [27, 28, 29, 30], which is the basis of the Frechet inception distance that is commonly used in machine learning [31]. If each two-dimensional distribution is a pile of sand, the EMD is the minimum amount of work required to transform one distribution into the other by moving sand around. EMD is easily converted to a score with range $(0,1)$ via exponentiation (Methods). To apply this score, the finite collections of datapoints in each distribution are binned into histograms. (Note, the algorithm used here, from [25], does not require smoothing the data by KDE, though that is an alternative approach.) In its native form, the EMD has the undesirable feature of being sensitive to the overall number of data points; to avoid this scale-dependence, the real and synthetic data are generally normalized, as we do here, before computing the score.
We calculated the earth-mover’s score for each fit in Fig. 2 and found this score also exhibits grade inflation (Table I). Interestingly, it resulted in high scores in all cases where the training data exhibited sharp boundaries, regardless of whether the fit was of low quality, as in Figs. 2b and d, with scores of 0.982 and 0.959, respectively, or high quality, as in Fig. 2e, which it gave a score of 0.949. The earth-mover’s score did, however, result in (appropriately) low scores in the low-quality fits of the two multimodal distributions, Figs. 2a and c, with scores of 0.124 and 0.379, respectively. Of note, scaling the scores by choosing a different value of $k$ does not resolve the grade inflation in Figs. 2b and d, because scaling so that lower quality fits have lower scores has the adverse effect of also reducing the score of high-quality fits: because changing $k$ does not change the ordering of scores, $k$ cannot thread this needle. Also of note, the earth-mover’s score can be highly sensitive to outliers, because the optimal transport might be such that the outliers have to be moved over further distances compared to more central points (see Discussion).
TABLE I: Scores for the fits and samples in Fig. 2. Fig. 2a-d are considered very low- or low-quality fits deserving lower scores; Fig. 2e is considered a high-quality fit deserving higher scores. See Results for details. Standard deviations are reported for scores whose calculation involves a stochastic element (Methods); they are not for repeat sampling (see Confidence Intervals).